# Probing ALP Lepton Flavour Violation at $\mu$TRISTAN

### v1

@@ 2 ALP-lepton interactions
$\mathcal{L}_{a\ell\ell}\supset\sum_{\ell_{i}}\frac{\partial_{\mu}a}{2f_{a}}\,\bar{\ell_{i}}C_{\ell_{i}\ell_{i}}^{A}\gamma^{\mu}\gamma_{5}\ell_{i}+\sum_{i\neq j}\frac{\partial_{\mu}a}{2f_{a}}\,\bar{\ell_{i}}\gamma^{\mu}\left(C_{\ell_{i}\ell_{j}}^{V}+C_{\ell_{i}\ell_{j}}^{A}\gamma_{5}\right)\ell_{j}\,,$ (1)
@@
$\mathcal{L}_{a\ell\ell}\supset-\mathrm{i}\sum_{i,j}\left[\frac{C_{\ell_{i}\ell_{j}}^{V}(m_{\ell_{j}}-m_{\ell_{i}})}{{2}f_{a}}\,a\bar{\ell_{i}}\ell_{j}+\frac{C_{\ell_{i}\ell_{j}}^{A}(m_{\ell_{i}}+m_{\ell_{j}})}{{2}f_{a}}\,a\bar{\ell_{i}}\gamma_{5}\ell_{j}\right]\,.$ (2)
@@
$g_{ij}^{V}\equiv\frac{C_{\ell_{i}\ell_{j}}^{V}(m_{\ell_{j}}-m_{\ell_{i}})}{2f_{a}},\quad g_{ij}^{A}\equiv\frac{C_{\ell_{i}\ell_{j}}^{A}(m_{\ell_{j}}+m_{\ell_{i}})}{2f_{a}},$ (3)
@@
$g_{ii}^{A}=\frac{C_{\ell_{i}\ell_{i}}^{A}m_{\ell_{i}}}{f_{a}},\quad g_{ij}^{A,V}\simeq\frac{C_{\ell_{i}\ell_{j}}^{A,V}m_{\ell_{j}}}{2f_{a}}~{}~{}~{}(i<j)\,.$ (4)
@@
Throughout the paper, we do not consider ALP fields that are genuine massless Nambu-Goldstone boson. Instead, we introduce a mass term $\frac{1}{2}m_{a}a^{2}$, assuming that it stems from explicit $\text{U}(1)$ breaking, and we treat the ALP mass $m_{a}$ as a free parameter. Following from the above interactions, it is then clear that, if kinematics allows, the ALP will decay into lepton pairs. The resulting width reads [30]:
@@
where we have summed over the final state charges. As we will see in the following sections, we expect $\mu$TRISTAN to be able to probe LFV ALP couplings as small as $g^{A,V}_{ij}\sim 10^{-3}-10^{-4}$. Hence, Eq. (6) implies that ALPs whose decays into charged leptons are allowed will always decay promptly inside the detector for the range of couplings we are interested in. Therefore, in this case, ALPs can be searched through distinctive nearly-background-free LFV signals.
@@
Figure 1: Contours in the plane $m_{a}-g_{ij}^{A,V}~{}(=m_{\ell_{j}}\,C_{\ell_{i}\ell_{j}}^{A,V}/f_{a})$ of the proper decay length $c\tau_{a}$ of an ALP whose only non-vanishing couplings are $C_{e\mu}^{V}=C_{e\mu}^{A}$ (left panel) and an ALP with all couplings $C_{\ell_{i}\ell_{j}}^{A,V}$ ($i\neq j$) and $C_{\ell_{i}\ell_{i}}^{A}$ equal (right panel).
@@
We now illustrate the above discussion showing the ALP proper decay length for different scenarios. In the left panel of Figure 1, the ALP has only LFV interactions to electrons and muons with $C_{e\mu}^{V}=C_{e\mu}^{A}$. In this case, the ALP is always long-lived below the $a\to e\mu$ kinematic threshold, always short-lived above it unless the coupling $g^{A,V}_{e\mu}$ is smaller than $10^{-5}-10^{-6}$, values beyond the $\mu$TRISTAN reach, as we will see.
@@ 3.1 ALP-mediated Lepton Flavour Violation
Notice that we set $f_{a}=100$ GeV and the relevant couplings $C^{A,V}_{\ell_{i}\ell_{j}}=1$, which corresponds to $g^{A,V}_{e\mu}=5.3\times 10^{-4}$ and $g^{A,V}_{\ell\tau}=8.9\times 10^{-3}$, as one can see from Eq. (4). As we mentioned, the scale $f_{a}$ is related to the vacuum expectation value (vev) of the scalar field that breaks the underlying $\text{U}(1)$ symmetry. As a consequence, we expect that viable UV completions of the effective Lagrangian in Eq. (1) will require $f_{a}\gtrsim 100$ GeV, hence $g^{A,V}_{e\mu}\lesssim 10^{-3}$ and $g^{A,V}_{\ell\tau}\lesssim 10^{-2}$. Hence, the cross sections in Table 1 are arguably the largest possible within realistic ALP models and we can conclude that $\mu$TRISTAN would not be capable to test LFV ALPs through these processes.
@@ 4 $\mu$TRISTAN sensitivity to LFV ALPs
In the following, we are showing the expected 95% confidence level (CL) exclusion potential of $\mu$TRISTAN, corresponding to $\mathcal{S}=2$ in Eq. (7), assuming that an integrated luminosity of $\mathcal{L}=10~{}\text{ab}^{-1}$ can be achieved. However, the most constraining searches being practically free of SM background, one can easily rescale the limits on the couplings we present as $\sim\left(10~{}\text{ab}^{-1}/\mathcal{L}\right)^{1/2}$.
@@ 4.1 ALP production through $e^{-}\mu^{+}\to a\gamma$
Figure 2: Distribution of the photon transverse momentum, $p_{T}^{\gamma}$, for $e^{-}\mu^{+}\to\gamma+$ invisible events before cuts. The blue line denotes the signal from production of a light invisible ALP, with $C^{A}_{e\mu}/f_{a}=C^{V}_{e\mu}/f_{a}=0.01~{}\text{GeV}^{-1}$, corresponding to $g^{A}_{e\mu}\simeq g^{V}_{e\mu}\simeq 5.3\times 10^{-4}$. The red line is the SM background $e^{-}\mu^{+}\to\gamma\nu_{e}\bar{\nu}_{\mu}$. The number of events is calculated including both production cross sections and assuming for the integrated luminosity $\mathcal{L}=10~{}\text{ab}^{-1}$.
@@
ALP couplings $m_{a}$ (GeV) Signature (BR) SM background Cut $\sigma_{\text{B}}(\rm{fb})\times\varepsilon_{\rm B}$ $\varepsilon_{\text{S}}=\varepsilon_{\rm sel}\times\varepsilon_{\rm cut}$ $2\sigma$ limit on $g_{e\mu}^{V,A}$
@@
Table 3: Summary of the possible searches for ALPs produced in $e^{-}\mu^{+}\to a\gamma$ for different ALP masses and couplings. The upper block shows the results for exclusively LFV ALPs, the lower block for ALPs also with LFC couplings. The third column reports the branching ratio (BR) of the ALP decay under consideration — the process denoted by has $\text{BR}\simeq 100\%$ but only about 40% of the decays occur in the inner detector for the chosen value of $m_{a}$. The applied cuts (if any), the BG cross section $\sigma_{\text{B}}$ and efficiency $\varepsilon_{\text{B}}$, the signal efficiency $\varepsilon_{\text{S}}\equiv\varepsilon_{\text{sel}}\times\varepsilon_{\text{cut}}$ (see the main text for details) are shown in the fifth, sixth and seventh columns, respectively. In the last column, we report the resulting $2\sigma$ limit on $g_{e\mu}^{V,A}\simeq m_{\mu}C_{e\mu}^{V,A}/2f_{a}$ (with $g_{e\mu}^{V}=g_{e\mu}^{A}$) for $\mathcal{L}=10~{}\text{ab}^{-1}$. For the BG-free searches these limits scale as $\sim(10~{}\text{ab}^{-1}/\mathcal{L})^{1/2}$.
@@
In Table 3, we summarise signals and backgrounds for the different ALP masses and coupling scenarios discussed above. We also show the kinematic cuts that effectively reduce the backgrounds in case they are present, and then compute the background and signal cut efficiencies. In order to estimate the signal significance, we additionally include the effect of the particle identification capabaility and geometric acceptance of the detector, $\varepsilon_{\rm sel}$, assuming as a requirement for the pseudorapidity of photons and leptons $|\eta|<2.5$. We finally obtain a $2\sigma$ lower bound on the flavour violating coupling $g_{e\mu}^{V,A}$ for an integrated luminosity of 10 ab-1 using Poisson statistics as in Eq. (7). We find that $\mu$TRISTAN has the capability to constrain LFV ALP couplings as low as $g_{e\mu}^{V,A}\sim 10^{-3}-10^{-4}$.
@@ 4.2 ALP production through LFV $\tau$ interactions
Figure 6: Left panel: total cross section of ALP production processes through LFV ALP-$\tau$ interactions at $\mu$TRISTAN for $|\eta|_{\text{max}}=2.5$ (solid lines) and $|\eta|_{\text{max}}=3.5$ (dashed lines), for $C^{A}_{\ell\tau}/f_{a}=C^{V}_{\ell\tau}/f_{a}=0.01~{}\text{GeV}^{-1}$ ($\ell=e\,\text{or}\,\mu$), corresponding to $g^{A}_{\ell\tau}\simeq g^{V}_{\ell\tau}\simeq 8.9\times 10^{-3}$. Right panel: signal efficiency of searches for the processes of the left plot with hadronic taus in the final state.
@@ 5 Summary and conclusions
  • For each signal channel, the analysis is divided into two mass regimes for the ALP and limits are set on dimensionless coupling $g_{ij}^{V,A}\simeq m_{j}C_{ij}^{V,A}/2f_{a}$ at an integrated luminosity of $\mathcal{L}=10$ ab-1. The limits scale as $\sim\left(10~{}\text{ab}^{-1}/\mathcal{L}\right)^{1/2}$ for background free signals (that are usually the most constraining ones).
@@
    • The search for heavy ALP through $e^{-}\mu^{+}\to a\gamma$ can test the parameter space beyond the limits from $\overline{\text{Mu}}$ oscillation and $\mu\to e\gamma$ and has the potential to test the $g^{V,A}_{e\mu}$ couplings to values as low as $\sim\mathcal{O}(10^{-4})$.
    • For the associated productions of ALP–$\tau$$\ell$ with $m_{a}>m_{\tau}+m_{\ell}$, the leptonic (hadronic) $\tau$ decay mode is sensitive to relatively low (high) ALP mass values. Signatures with hadronic $\tau$ decays are able to probe values of the $g_{\ell\tau}^{V,A}$ couplings currently allowed by $\tau$ LFV decay and leptonic $g-2$ and reach a sensitivity of $g_{\mu\tau}^{V,A}\sim 10^{-3}-10^{-2}$ and $g^{V,A}_{e\tau}\sim 10^{-2}$ for $m_{a}=\mathcal{O}(100)$ GeV.
@@
As a future outlook, one can envisage different muon beam energy configurations for the $\mu$TRISTAN project. We have seen that the ALP production cross-section for the $e^{-}\mu^{+}$ collisions increases for lower c.m. energies, cf. Eq. (9). Furthermore, if the asymmetry in the beam energies is lowered, the detecting efficiency of the signal also increases, as the ALP decay products are less collinear to the muon beam. We find that the optimal choice for the beam energy configuration is $(E_{e},E_{\mu})=(30~{}\rm{GeV},100~{}\rm{GeV})$, [^6] Such a configuration corresponds to $\sqrt{s}\simeq 110$ GeV, hence it is clearly beyond the main purpose of $\mu$TRISTAN to work as a Higgs factory. However, if evidence of light NP in the $\sqrt{s}\simeq 346$ GeV run is observed, performing precision studies of it at lower $\sqrt{s}$ is a reasonable option. and study how this influences the collider bound on $g_{e\mu}^{V,A}$ that can be obtained in the search for $e^{-}\mu^{+}\to\gamma a(\to e^{+}\mu^{-}/\mu^{+}\mu^{-})$. Although in such a case the luminosity is expected to be lower due to relativistic effects — for more details see [64] — we show in Figure 9 the improvement on the sensitivity to $g_{e\mu}^{V,A}$ (by approximately one order of magnitude) assuming that an integrated luminosity of $10~{}\rm{ab}^{-1}$ can still be reached in the future with advanced beam intensity and focusing technologies. For the ALP production modes with an associated $\tau$ lepton in the final state, we observe that $\mu$TRISTAN with the above energy configuration does not exhibit a better performance compared to the results obtained in this paper, although the cross sections also increase, since the $\tau$-jet identification efficiency rapidly deteriorates for softer jets.

### v2

@@ 2 ALP-lepton interactions
$\mathcal{L}_{a\ell\ell}\supset\sum_{i}\frac{\partial_{\mu}a}{2f_{a}}\,\bar{\ell_{i}}C_{\ell_{i}\ell_{i}}^{A}\gamma^{\mu}\gamma_{5}\ell_{i}+\sum_{i\neq j}\frac{\partial_{\mu}a}{2f_{a}}\,\bar{\ell_{i}}\gamma^{\mu}\left(C_{\ell_{i}\ell_{j}}^{V}+C_{\ell_{i}\ell_{j}}^{A}\gamma_{5}\right)\ell_{j}\,,$ (1)
@@
$\mathcal{L}_{a\ell\ell}\supset-\mathrm{i}\sum_{i,j}\left[\frac{C_{\ell_{i}\ell_{j}}^{V}(m_{\ell_{i}}-m_{\ell_{j}})}{{2}f_{a}}\,a\bar{\ell_{i}}\ell_{j}+\frac{C_{\ell_{i}\ell_{j}}^{A}(m_{\ell_{i}}+m_{\ell_{j}})}{{2}f_{a}}\,a\bar{\ell_{i}}\gamma_{5}\ell_{j}\right]\,.$ (2)
@@
$g_{ij}^{V}\equiv\frac{C_{\ell_{i}\ell_{j}}^{V}(m_{\ell_{i}}-m_{\ell_{j}})}{2f_{a}},\quad g_{ij}^{A}\equiv\frac{C_{\ell_{i}\ell_{j}}^{A}(m_{\ell_{j}}+m_{\ell_{i}})}{2f_{a}},$ (3)
@@
$g_{ii}^{A}=\frac{C_{\ell_{i}\ell_{i}}^{A}m_{\ell_{i}}}{f_{a}},\quad g_{ij}^{A,V}\simeq\pm\frac{C_{\ell_{i}\ell_{j}}^{A,V}m_{\ell_{j}}}{2f_{a}}~{}~{}~{}(i<j)\,.$ (4)
@@
Throughout the paper, we do not consider ALP fields that are genuine massless Nambu-Goldstone boson. Instead, we introduce a mass term $\frac{1}{2}m_{a}^{2}a^{2}$, assuming that it stems from explicit $\text{U}(1)$ breaking, and we treat the ALP mass $m_{a}$ as a free parameter. Following from the above interactions, it is then clear that, if kinematics allows, the ALP will decay into lepton pairs. The resulting width reads [30]:
@@
where we have summed over the final state charges. As we will see in the following sections, we expect $\mu$TRISTAN to be able to probe LFV ALP couplings as small as $|g^{A,V}_{ij}|\sim 10^{-3}-10^{-4}$. Hence, Eq. (6) implies that ALPs whose decays into charged leptons are allowed will always decay promptly inside the detector for the range of couplings we are interested in. Therefore, in this case, ALPs can be searched through distinctive nearly-background-free LFV signals.
@@
Figure 1: Contours in the plane $m_{a}-|g_{ij}^{A,V}|~{}(=m_{\ell_{j}}\,|C_{\ell_{i}\ell_{j}}^{A,V}|/f_{a})$ of the proper decay length $c\tau_{a}$ of an ALP whose only non-vanishing couplings are $C_{e\mu}^{V}=C_{e\mu}^{A}$ (left panel) and an ALP with all couplings $C_{\ell_{i}\ell_{j}}^{A,V}$ ($i\neq j$) and $C_{\ell_{i}\ell_{i}}^{A}$ equal (right panel).
@@
We now illustrate the above discussion showing the ALP proper decay length for different scenarios. In the left panel of Figure 1, the ALP has only LFV interactions to electrons and muons with $C_{e\mu}^{V}=C_{e\mu}^{A}$. In this case, the ALP is always long-lived below the $a\to e\mu$ kinematic threshold, always short-lived above it unless the coupling $|g^{A,V}_{e\mu}|$ is smaller than $10^{-5}-10^{-6}$, values beyond the $\mu$TRISTAN reach, as we will see.
@@ 3.1 ALP-mediated Lepton Flavour Violation
Notice that we set $f_{a}=100$ GeV and the relevant couplings $C^{A,V}_{\ell_{i}\ell_{j}}=1$, which corresponds to $|g^{A,V}_{e\mu}|=5.3\times 10^{-4}$ and $|g^{A,V}_{\ell\tau}|=8.9\times 10^{-3}$, as one can see from Eq. (4). As we mentioned, the scale $f_{a}$ is related to the vacuum expectation value (vev) of the scalar field that breaks the underlying $\text{U}(1)$ symmetry. As a consequence, we expect that viable UV completions of the effective Lagrangian in Eq. (1) will require $f_{a}\gtrsim 100$ GeV, hence $|g^{A,V}_{e\mu}|\lesssim 10^{-3}$ and $|g^{A,V}_{\ell\tau}|\lesssim 10^{-2}$. Hence, the cross sections in Table 1 are arguably the largest possible within realistic ALP models and we can conclude that $\mu$TRISTAN would not be capable to test LFV ALPs through these processes.
@@ 4 $\mu$TRISTAN sensitivity to LFV ALPs
In the following, we are showing the expected exclusion potential of $\mu$TRISTAN at the 95% confidence level (CL), corresponding to $\mathcal{S}=2$ in Eq. (7), assuming that an integrated luminosity of $\mathcal{L}=10~{}\text{ab}^{-1}$ can be achieved. However, the most constraining searches being practically free of SM background, one can easily rescale the limits on the couplings we present as $\sim\left(10~{}\text{ab}^{-1}/\mathcal{L}\right)^{1/2}$.
@@ 4.1 ALP production through $e^{-}\mu^{+}\to a\gamma$
Figure 2: Distribution of the photon transverse momentum, $p_{T}^{\gamma}$, for $e^{-}\mu^{+}\to\gamma+$ invisible events before cuts. The blue line denotes the signal from production of a light invisible ALP, with $C^{A}_{e\mu}/f_{a}=C^{V}_{e\mu}/f_{a}=0.01~{}\text{GeV}^{-1}$, corresponding to $|g^{A}_{e\mu}|\simeq|g^{V}_{e\mu}|\simeq 5.3\times 10^{-4}$. The red line is the SM background $e^{-}\mu^{+}\to\gamma\nu_{e}\bar{\nu}_{\mu}$. The number of events is calculated including both production cross sections and assuming for the integrated luminosity $\mathcal{L}=10~{}\text{ab}^{-1}$.
@@
ALP couplings $m_{a}$ (GeV) Signature (BR) SM background Cut $\sigma_{\text{B}}(\rm{fb})\times\varepsilon_{\rm B}$ $\varepsilon_{\text{S}}=\varepsilon_{\rm sel}\times\varepsilon_{\rm cut}$ $2\sigma$ limit on $|g_{e\mu}^{V,A}|$
@@
Table 3: Summary of the possible searches for ALPs produced in $e^{-}\mu^{+}\to a\gamma$ for different ALP masses and couplings. The upper block shows the results for exclusively LFV ALPs, the lower block for ALPs also with LFC couplings. The third column reports the branching ratio (BR) of the ALP decay under consideration — the process denoted by has $\text{BR}\simeq 100\%$ but only about 40% of the decays occur in the inner detector for the chosen value of $m_{a}$. The applied cuts (if any), the BG cross section $\sigma_{\text{B}}$ and efficiency $\varepsilon_{\text{B}}$, the signal efficiency $\varepsilon_{\text{S}}\equiv\varepsilon_{\text{sel}}\times\varepsilon_{\text{cut}}$ (see the main text for details) are shown in the fifth, sixth and seventh columns, respectively. In the last column, we report the resulting $2\sigma$ limit on $|g_{e\mu}^{V,A}|\simeq m_{\mu}|C_{e\mu}^{V,A}|/2f_{a}$ (with $C_{e\mu}^{V}=C_{e\mu}^{A}$) for $\mathcal{L}=10~{}\text{ab}^{-1}$. For the BG-free searches these limits scale as $\sim(10~{}\text{ab}^{-1}/\mathcal{L})^{1/2}$.
@@
In Table 3, we summarise signals and backgrounds for the different ALP masses and coupling scenarios discussed above. We also show the kinematic cuts that effectively reduce the backgrounds in case they are present, and then compute the background and signal cut efficiencies. In order to estimate the signal significance, we additionally include the effect of the particle identification capabaility and geometric acceptance of the detector, $\varepsilon_{\rm sel}$, assuming as a requirement for the pseudorapidity of photons and leptons $|\eta|<2.5$. We finally obtain a $2\sigma$ lower bound on the flavour violating coupling $g_{e\mu}^{V,A}$ for an integrated luminosity of 10 ab-1 using Poisson statistics as in Eq. (7). We find that $\mu$TRISTAN has the capability to constrain LFV ALP couplings as low as $|g_{e\mu}^{V,A}|\sim 10^{-3}-10^{-4}$.
@@ 4.2 ALP production through LFV $\tau$ interactions
Figure 6: Left panel: total cross section of ALP production processes through LFV ALP-$\tau$ interactions at $\mu$TRISTAN for $|\eta|_{\text{max}}=2.5$ (solid lines) and $|\eta|_{\text{max}}=3.5$ (dashed lines), for $C^{A}_{\ell\tau}/f_{a}=C^{V}_{\ell\tau}/f_{a}=0.01~{}\text{GeV}^{-1}$ ($\ell=e\,\text{or}\,\mu$), corresponding to $|g^{A}_{\ell\tau}|\simeq|g^{V}_{\ell\tau}|\simeq 8.9\times 10^{-3}$. Right panel: signal efficiency of searches for the processes of the left plot with hadronic taus in the final state.
@@ 5 Summary and conclusions
  • For each signal channel, the analysis is divided into two mass regimes for the ALP and limits are set on dimensionless coupling $|g_{ij}^{V,A}|\simeq m_{j}C_{ij}^{V,A}/2f_{a}$ at an integrated luminosity of $\mathcal{L}=10$ ab-1. The limits scale as $\sim\left(10~{}\text{ab}^{-1}/\mathcal{L}\right)^{1/2}$ for background free signals (that are usually the most constraining ones).
@@
    • The search for heavy ALP through $e^{-}\mu^{+}\to a\gamma$ can test the parameter space beyond the limits from $\overline{\text{Mu}}$ oscillation and $\mu\to e\gamma$ and has the potential to test the $|g^{V,A}_{e\mu}|$ couplings to values as low as $\sim\mathcal{O}(10^{-4})$.
    • For the associated productions of ALP–$\tau$$\ell$ with $m_{a}>m_{\tau}+m_{\ell}$, the leptonic (hadronic) $\tau$ decay mode is sensitive to relatively low (high) ALP mass values. Signatures with hadronic $\tau$ decays are able to probe values of the $g_{\ell\tau}^{V,A}$ couplings currently allowed by $\tau$ LFV decay and leptonic $g-2$ and reach a sensitivity of $|g_{\mu\tau}^{V,A}|\sim 10^{-3}-10^{-2}$ and $|g^{V,A}_{e\tau}|\sim 10^{-2}$ for $m_{a}=\mathcal{O}(100)$ GeV.
The $\mu$TRISTAN sensitivity reach summarised above approximately corresponds to the following limits on the cut-off scale of the relevant operators in Eq. (1):
${f_{a}}/{C^{V,A}_{e\mu}}\lesssim 500~{}\text{GeV}\,,\quad{f_{a}}/{C^{V,A}_{\mu\tau}}\lesssim 300~{}\text{GeV}\,,\quad{f_{a}}/{C^{V,A}_{e\tau}}\lesssim 150~{}\text{GeV}\,.$ (10)
Comparing these results with the $\mu$TRISTAN $e^{-}\mu^{+}$ mode centre of mass energy, $\sqrt{s}\simeq 346$ GeV, we can conclude that our model-independent effective field theory approach is marginally justified in the case of searches for $e$-$\mu$ violation. On the other hand, the $\mu$TRISTAN sensitivity on $\mu$-$\tau$ and, especially, $e$-$\tau$ ALP LFV searches would be within the validity of the effective field theory only if $C^{V,A}_{\ell\tau}>1$, which would point to an ALP originating from a strongly-coupled sector. The search for $\mu$-$\tau$ LFV with $\sqrt{s}=2$ TeV $\mu^{+}\mu^{+}$ collisions is definitely not sensitive to ranges of couplings for which our effective field theory is consistent.
It is a model-dependent question whether the fields associated with the UV completion substantially affect the ALP phenomenology. Likewise, LHC limits on the mass of new particles charged under the SM electroweak interactions depend on the specific production and decay modes and range approximately between $100$ GeV and $1$ TeV, hence it will depend on the UV completion whether $\mu$TRISTAN is capable to test ALP models beyond the LHC constraints or not. While this important discussion is beyond the scope of our work, we just notice here that a UV-complete ALP model might entail fields charged under the SM gauge group at a scale parametrically larger than $f_{a}$, which is the vev of the singlet scalar field that breaks the associated $\text{U}(1)$ symmetry. [^6] This is for instance the case of the “familon”, the PNGB arising from a Froggatt-Nielsen $\text{U}(1)$ flavour symmetry. The UV completion of such flavour models (e.g. consisting in vectorlike fermions) typically lie at a scale $\Lambda/f_{a}\gtrsim\mathcal{O}(10)$, see e.g. [30].
As a future outlook, one can envisage different muon beam energy configurations for the $\mu$TRISTAN project. We have seen that the ALP production cross-section for the $e^{-}\mu^{+}$ collisions increases for lower c.m. energies, cf. Eq. (9). Furthermore, if the asymmetry in the beam energies is lowered, the detecting efficiency of the signal also increases, as the ALP decay products are less collinear to the muon beam. We find that the optimal choice for the beam energy configuration is $(E_{e},E_{\mu})=(30~{}\rm{GeV},100~{}\rm{GeV})$, [^7] Such a configuration corresponds to $\sqrt{s}\simeq 110$ GeV, hence it is clearly beyond the main purpose of $\mu$TRISTAN to work as a Higgs factory. However, if evidence of light NP in the $\sqrt{s}\simeq 346$ GeV run is observed, performing precision studies of it at lower $\sqrt{s}$ is a reasonable option. and study how this influences the collider bound on $g_{e\mu}^{V,A}$ that can be obtained in the search for $e^{-}\mu^{+}\to\gamma a(\to e^{+}\mu^{-}/\mu^{+}\mu^{-})$. Although in such a case the luminosity is expected to be lower due to relativistic effects — for more details see [64] — we show in Figure 9 the improvement on the sensitivity to $g_{e\mu}^{V,A}$ (by approximately one order of magnitude) assuming that an integrated luminosity of $10~{}\rm{ab}^{-1}$ can still be reached in the future with advanced beam intensity and focusing technologies. We notice that this improved sensitivity corresponds to ${f_{a}}/{C^{V,A}_{e\mu}}$ above the TeV scale. In combination with a lower value of $\sqrt{s}$, this enlarges the validity range of the the effective Lagrangian in Eq. (1). For the ALP production modes with an associated $\tau$ lepton in the final state, we observe that $\mu$TRISTAN with the above energy configuration does not exhibit a better performance compared to the results obtained in this paper, although the cross sections also increase, since the $\tau$-jet identification efficiency rapidly deteriorates for softer jets.